# Neural Approximate Mirror Maps for Constrained Diffusion Models

### v1

@@ Abstract
Diffusion models excel at creating visually-convincing images, but they often struggle to meet subtle constraints inherent in the training data. Such constraints could be physics-based (e.g., satisfying a PDE), geometric (e.g., respecting symmetry), or semantic (e.g., including a particular number of objects). When the training data all satisfy a certain constraint, enforcing this constraint on a diffusion model not only improves its distribution-matching accuracy but also makes it more reliable for generating valid synthetic data and solving constrained inverse problems. However, existing methods for constrained diffusion models are inflexible with different types of constraints. Recent work proposed to learn mirror diffusion models (MDMs) in an unconstrained space defined by a mirror map and to impose the constraint with an inverse mirror map, but analytical mirror maps are challenging to derive for complex constraints. We propose neural approximate mirror maps (NAMMs) for general constraints. Our approach only requires a differentiable distance function from the constraint set. We learn an approximate mirror map that pushes data into an unconstrained space and a corresponding approximate inverse that maps data back to the constraint set. A generative model, such as an MDM, can then be trained in the learned mirror space and its samples restored to the constraint set by the inverse map. We validate our approach on a variety of constraints, showing that compared to an unconstrained diffusion model, a NAMM-based MDM substantially improves constraint satisfaction. We also demonstrate how existing diffusion-based inverse-problem solvers can be easily applied in the learned mirror space to solve constrained inverse problems.
@@ 1 Introduction
Figure 1: Conceptual illustration. (a) Despite being trained on a data distribution constrained to $\mathcal{M}$, a regular diffusion model (DM) may generate samples that violate the constraint due to learning and/or sampling error. (b) We propose to learn a neural approximate mirror map (NAMM) that entails both a forward map $\mathbf{g}_{\phi}$ and inverse map $\mathbf{f}_{\psi}$. The forward map transforms the constrained space into an unconstrained (“mirror”) space. Once $\mathbf{g}_{\phi}$ and $\mathbf{f}_{\psi}$ are learned, a mirror diffusion model (MDM) can be trained on the pushforward of the data distribution through $\mathbf{g}_{\phi}$. The approximate inverse map $\mathbf{f}_{\psi}$ maps MDM samples back to the constrained space.
@@
Many data distributions follow a rule that is not visually obvious. For example, videos of fluid flow obey a partial differential equation (PDE), but a human may find it difficult to discern whether a given video truly agrees with the prescribed PDE. We can characterize such distributions as constrained distributions. Theoretically, a generative model trained on a constrained image distribution should satisfy the constraint, but in practice, it may generate visually-convincing images that break the rules. Constraint violation is caused by learning error and sampling error [20]. Ensuring constraint satisfaction in spite of learning and sampling errors would make generative models more reliable for applications such as solving inverse problems with physics-informed deep-learned priors.
@@
Diffusion models in particular are popular types of generative models, but existing approaches for incorporating constraints either restrict the type of constraint or do not scale well. Equivariant score-based models [51], Riemannian diffusion models [21, 36], reflected diffusion models [48, 26], log-barrier methods [26], and mirror diffusion models (as originally proposed) [47] are restricted to certain types of constraints, such as symmetry groups [51], Riemannian manifolds [21, 36], or convex sets [26, 47]. Generally speaking, these restrictions are in the service of guaranteeing a hard constraint, whereas a soft constraint offers flexibility and may be sufficient in many scenarios. One could, for example, introduce a guidance term to encourage constraint satisfaction during sampling [32, 6, 71], but so far there does not exist a principled framework to do so. Previous work suggested imposing a soft constraint during training [20] by estimating the clean image at every noisy diffusion step and evaluating its constraint satisfaction. However, approximation [23] of the clean image is extremely crude at high noise levels and thus unsuitable for constraints that are sensitive to approximation error and noise, such as a PDE constraint. Finally, it is always possible to impose the constraint on samples generated by the diffusion model, such as by rejecting invalid samples, either during or after training [37]. However, these approaches rely on computationally-expensive simulation steps. Instead, we aim for a flexible approach to impose general constraints by construction.
@@
Our goal is to find an invertible function that maps constrained images into an unconstrained space so that a regular generative model can be trained in the unconstrained space and automatically satisfy the constraint through the inverse function. We propose neural approximate mirror maps (NAMMs), which bring the flexibility of soft constraints into the principled framework of mirror diffusion models [47]. A mirror diffusion model (MDM) allows for training a completely unconstrained diffusion model in a “mirror” space defined by a mirror map. Unconstrained samples from the trained diffusion model are mapped back to the constrained space via an inverse mirror map. However, invertible mirror maps are challenging to derive in closed form for general (e.g., non-convex) constraints.
@@
A NAMM encompasses a (forward) mirror map $\mathbf{g}_{\phi}$ and its approximate inverse map $\mathbf{f}_{\psi}$, which are trained to offer the desired benefits of an analytical mirror map. In particular, $\mathbf{f}_{\psi}\approx\mathbf{g}_{\phi}^{-1}$, and $\mathbf{f}_{\psi}$ maps unconstrained points to the constrained space (see Fig. 1 for a conceptual illustration). Our method works for any constraint that has a differentiable function to quantify the distance from an image to the constraint set. We parameterize $\mathbf{g}_{\phi}$ as the gradient of a strongly input-convex neural network (ICNN) [4] to satisfy invertibility. We train the NAMM with a cycle-consistency loss [73] to ensure $\mathbf{f}_{\psi}\approx{\mathbf{g}_{\phi}}^{-1}$ and train the inverse map with a constraint loss to ensure $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})$ is close to the constraint set for all $\tilde{\mathbf{x}}$ that we are interested in (we will define this more formally). An MDM can be trained on the pushforward of the data distribution through $\mathbf{g}_{\phi}$, and its generated samples can be mapped to the constraint set via $\mathbf{f}_{\psi}$. Although not inherently restricted to diffusion models, our approach maintains the many advantages of diffusion models, including expressive generation, simulation-free training [62], and tractable computation of probability densities [47]. One can also adapt existing diffusion-based inverse solvers for the mirror space and enforce constraints with the inverse map.
@@
Through our experiments, we show improved constraint satisfaction for various physics-based, geometric, and semantic constraints. We demonstrate how a NAMM can be used to solve constrained inverse problems, such as data assimilation. Finally, we validate the soundness of the method through several ablation studies, in particular finding that modeling the mirror map as the gradient of a strongly-convex function brings not just theoretical, but also practical, benefits.
@@ 2.1 Constrained generative models
Explicitly incorporating a data constraint into a generative model poses benefits such as data efficiency [28, 8], generalization capabilities [45], and feasibility of samples [31]. Some methods leverage equivariant neural networks [57, 30, 66] for symmetry [3, 51, 43, 64, 11, 55, 29, 45, 50, 22, 69, 34, 70] but do not generalize to other types of constraints and are often restricted to a certain type of generative model (e.g., flow-based or score-based models) [43, 64, 11, 55, 29, 45, 50, 51, 69, 34, 70, 22, 3]. Recent work has studied constrained diffusion models, including Riemannian diffusion models [21, 36], log-barrier methods [26], reflected diffusion models [48], and mirror diffusion models [47]. However, these methods make strong assumptions about the constraint, such as being characterized as a Riemannian manifold [21, 36], having a well-defined reflection operator [48], or corresponding to a convex constraint set [26, 47]. Fishman et al. [27] proposed a diffusion model that incorporates Metropolis-Hastings steps to be compatible with a general constraint, but impractically high rejection rates may occur with constraints that are challenging to satisfy, such as a PDE constraint.
An alternative approach is to introduce a soft constraint penalty when training the generative model [28, 49, 15, 31, 20]. However, evaluating the constraint loss of generated samples during training may be prohibitively expensive. Instead, one could add constraint-violating training examples [31], but it is often difficult to procure useful negative examples. In contrast to previous soft approaches, our approach does not alter the training objective of the generative model.
@@ 2.2 Mirror maps
For any convex constraint set $\mathcal{C}\subseteq\mathbb{R}^{d}$, one can define a mirror map that maps from the constraint set to all of $\mathbb{R}^{d}$. This is done by first defining a mirror potential $\phi:\mathcal{C}\to\mathbb{R}$ that is continuously-differentiable and strongly-convex [12, 65]. The mirror map is defined as the gradient $\nabla\phi:\mathcal{C}\to\mathbb{R}^{d}$ [47]. Importantly, every mirror map has an inverse $(\nabla\phi)^{-1}:\mathbb{R}^{d}\to\mathcal{C}$. Unlike the forward mirror map, the inverse map is not necessarily the gradient of a strongly-convex function [72, 65]. Mirror maps have been used for constrained optimization [9] and constrained sampling [35, 46, 47].
Although true mirror maps exist only for convex constraint sets, we seek to generalize the concept to learn approximate mirror maps that handle more complex constraint sets. Recent work suggested learning mirror maps for purposes such as convex optimization [65] and reinforcement learning [2] but did not tackle constrained generative modeling.
@@ 2.3 Diffusion models
A diffusion model is a type of generative model that learns to generate new data samples through a gradual denoising process [58, 33, 60, 62, 42]. The diffusion, or noising, process can be modeled as a stochastic differential equation (SDE) [62]. The diffusion SDE induces a time-dependent image distribution $p_{t}$, where $p_{0}=p_{\text{data}}$ (the target image distribution) and $p_{T}\approx\mathcal{N}(\mathbf{0},\mathbf{I})$. The goal of a diffusion model is to generate new images by reversing the diffusion process. This can be done with the score function of $p_{t}$, defined as $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. In the context of imaging problems, the score function is often parameterized using a convolutional neural network (CNN) with parameters $\theta$, which we denote by $\mathbf{s}_{\theta}$. The score model is trained with a “denoising score matching” objective [38, 67, 62] that allows the time-dependent score to be learned in a simulation-free manner (i.e., it does not require simulating the forward noising process during training). The score function appears in a reverse-time SDE [62, 5] that can be used to sample from the clean image distribution $p_{0}$ by first sampling an image of pure noise from $\mathcal{N}(\mathbf{0},\mathbf{I})$ and then gradually denoising it.
@@ 3 Method
We will now describe neural approximate mirror maps (NAMMs) for constrained diffusion models. While we focus on diffusion models, any generative model can be trained in the learned mirror space. We denote the constrained image distribution by $p_{\text{data}}$ and the (not necessarily convex) constraint set by $\mathcal{M}\subseteq\mathbb{R}^{d}$. Images in the constrained space and mirror space are denoted by $\mathbf{x}$ and $\tilde{\mathbf{x}}$, respectively.
@@ 3.1 Learning the forward and inverse mirror maps
Figure 2: NAMM training illustration. Given data that lie on a constraint manifold $\mathcal{M}$ (e.g., the hyperplane of images with the same total brightness), we jointly train an approximate mirror map $\mathbf{g}_{\phi}$ and its approximate inverse $\mathbf{f}_{\psi}$. After mapping data $\mathbf{x}\sim p_{\text{data}}$ to the mirror space as $\mathbf{g}_{\phi}(\mathbf{x})$, we perturb them with additive Gaussian noise whose std. dev. can be anywhere between 00 and $\sigma_{\text{max}}$. The inverse map $\mathbf{f}_{\psi}$ is trained to map these perturbed samples back onto $\mathcal{M}$.
@@
A true inverse mirror map satisfies cycle consistency and constraint satisfaction on all of $\mathbb{R}^{d}$, so ideally $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\mathbf{g}_{\phi}^{-1}(\tilde{\mathbf{x}})$ and $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})\in\mathcal{M}$ for all $\tilde{\mathbf{x}}\in\mathbb{R}^{d}$. But since it would be computationally infeasible to optimize $\mathbf{f}_{\psi}$ over all possible points in $\mathbb{R}^{d}$, we instead optimize it over distributions that we would expect the inverse map to face in practice in the context of diffusion models. That is, we only need $\mathbf{f}_{\psi}$ to be valid for samples from an MDM trained on ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$, which we refer to as the mirror distribution. To make $\mathbf{f}_{\psi}$ robust to learning/sampling error of the MDM, we consider a sequence of noisy distributions in the mirror space, each corresponding to adding Gaussian noise to samples from ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$, which, given a maximum perturbation level $\sigma_{\text{max}}$, we denote by
@@
We train $\mathbf{f}_{\psi}$ to be a valid inverse mirror map only for points from these noisy mirror distributions. Since we do not know a priori how much noise the MDM samples will contain, we consider all possible noise levels up to $\sigma_{\text{max}}$ for robustness (see Appendix A.2 for an ablation study of this choice).
We define the following cycle-consistency loss [73], which covers both the forward and inverse directions and evaluates the inverse direction for the entire sequence of distributions defined in Eq. 2:
@@
Let $\ell_{\text{constr}}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be a differentiable constraint distance that measures the distance from an input image to the constraint set. We define the following constraint loss to encourage $\mathbf{f}_{\psi}$ to map points from the noisy mirror distributions (Eq. 2) to the constraint set:
@@
We use Monte-Carlo to approximate the expectations in the objective over the noisy mirror distributions with $\sigma\sim\mathcal{U}([0,\sigma_{\text{max}}])$ and approximately solve Eq. 1 with stochastic gradient descent.
@@
We parameterize $\mathbf{g}_{\phi}$ as the gradient of an input-convex neural network (ICNN) following the implementation of Tan et al. [65], thus satisfying the requirement of being the gradient of a strongly-convex function. We note that $\mathbf{g}_{\phi}$ is not a true mirror map since $\mathcal{M}$ is not assumed to be convex, and $\mathbf{g}_{\phi}$ is defined on all of $\mathbb{R}^{d}$ instead of just on $\mathcal{M}$. We parameterize $\mathbf{f}_{\psi}$ as a ResNet-based CNN similar to the one used in CycleGAN [73]. Note that $\mathbf{f}_{\psi}$ is also the gradient of a strongly-convex function if the cycle-consistency loss is 00 (i.e., $\mathbf{f}_{\psi}=\mathbf{g}_{\phi}^{-1}$).

### v2

@@ Abstract
Diffusion models excel at creating visually-convincing images, but they often struggle to meet subtle constraints inherent in the training data. Such constraints could be physics-based (e.g., satisfying a PDE), geometric (e.g., respecting symmetry), or semantic (e.g., including a particular number of objects). When the training data all satisfy a certain constraint, enforcing this constraint on a diffusion model makes it more reliable for generating valid synthetic data and solving constrained inverse problems. However, existing methods for constrained diffusion models are restricted in the constraints they can handle. For instance, recent work proposed to learn mirror diffusion models (MDMs), but analytical mirror maps only exist for convex constraints and can be challenging to derive. We propose neural approximate mirror maps (NAMMs) for general, possibly non-convex constraints. Our approach only requires a differentiable distance function from the constraint set. We learn an approximate mirror map that transforms data into an unconstrained space and a corresponding approximate inverse that maps data back to the constraint set. A generative model, such as an MDM, can then be trained in the learned mirror space and its samples restored to the constraint set by the inverse map. We validate our approach on a variety of constraints, showing that compared to an unconstrained diffusion model, a NAMM-based MDM substantially improves constraint satisfaction. We also demonstrate how existing diffusion-based inverse-problem solvers can be easily applied in the learned mirror space to solve constrained inverse problems.
@@ 1 Introduction
Many data distributions follow a rule that is not visually obvious. For example, videos of fluid flow obey a partial differential equation (PDE), but a human may find it difficult to discern whether a video agrees with the prescribed PDE. We can characterize such distributions as constrained distributions. Theoretically, a generative model trained on a constrained image distribution should satisfy the constraint, but in practice due to learning and sampling errors (Daras et al., 2024), it may generate visually-convincing images that break the rules. Ensuring constraint satisfaction in spite of such errors would make generative models more reliable for applications such as solving inverse problems.
Figure 1: Conceptual illustration. (a) Despite being trained on a data distribution constrained to $\mathcal{M}$, a regular diffusion model (DM) may generate samples that violate the constraint. (b) We propose to learn a neural approximate mirror map (NAMM) that entails a forward map $\mathbf{g}_{\phi}$ and inverse map $\mathbf{f}_{\psi}$. The forward map transforms the constrained space into an unconstrained (“mirror”) space. Once $\mathbf{g}_{\phi}$ and $\mathbf{f}_{\psi}$ are learned, a mirror diffusion model (MDM) can be trained on the pushforward of the data distribution through $\mathbf{g}_{\phi}$ and its samples mapped back to the constrained space through $\mathbf{f}_{\psi}$.
@@
Diffusion models are popular generative models, but existing approaches for incorporating constraints either restrict the type of constraint or do not scale well. Equivariant (Niu et al., 2020), Riemannian (De Bortoli et al., 2022; Huang et al., 2022), reflected (Lou & Ermon, 2023; Fishman et al., 2023a), log-barrier (Fishman et al., 2023a), and mirror (Liu et al., 2024) diffusion models are all restricted to certain types of constraints, such as symmetry groups (Niu et al., 2020), Riemannian manifolds (De Bortoli et al., 2022; Huang et al., 2022), or convex sets (Fishman et al., 2023a; Liu et al., 2024). Generally speaking, these restrictions are in the service of guaranteeing a hard constraint, whereas a soft constraint offers flexibility and may be sufficient in many scenarios. One could, for example, introduce a guidance term to encourage constraint satisfaction during sampling (Graikos et al., 2022; Bansal et al., 2023; Zhang et al., 2023), but so far there does not exist a principled framework to do so. Previous work suggested imposing a soft constraint during training (Daras et al., 2024; Upadhyay et al., 2023) by estimating the clean image at every noisy diffusion step and evaluating its constraint satisfaction. However, approximation (Efron, 2011) of the clean image from the intermediate noisy image is crude at high noise levels and thus unsuitable for constraints that are sensitive to approximation error and noise, such as a PDE constraint. Of course it is possible to penalize invalid generated samples during or after training (Huang et al., 2024), but this approach relies on computationally-expensive simulation steps. Instead, we aim for a flexible approach to impose general constraints by construction.
@@
Our goal is to find an invertible function that maps constrained images into an unconstrained space so that a regular generative model can be trained in the unconstrained space and automatically satisfy the constraint through the inverse function. We propose neural approximate mirror maps (NAMMs) [^1] Code can be found at https://github.com/berthyf96/namm., which bring the flexibility of soft constraints into the principled framework of mirror diffusion models (Liu et al., 2024). A mirror diffusion model (MDM) allows for training a completely unconstrained diffusion model in a “mirror” space defined by a mirror map. Unconstrained samples from the diffusion model are mapped back to the constrained space via an inverse mirror map. However, invertible mirror maps are challenging or impossible to derive in closed form for general constraints. We address this by jointly optimizing two networks to approximate a mirror map and its inverse.
@@
A NAMM encompasses a (forward) mirror map $\mathbf{g}_{\phi}$ and its approximate inverse map $\mathbf{f}_{\psi}$. They are trained so that $\mathbf{f}_{\psi}\approx\mathbf{g}_{\phi}^{-1}$, and $\mathbf{f}_{\psi}$ maps unconstrained points to the constrained space (see Figure 1 for a conceptual illustration). Our method works for any constraint that has a differentiable function to quantify the distance from an image to the constraint set. We parameterize $\mathbf{g}_{\phi}$ as the gradient of a strongly input-convex neural network (ICNN) (Amos et al., 2017) to satisfy invertibility. We train the NAMM with a cycle-consistency loss (Zhu et al., 2017) to ensure $\mathbf{f}_{\psi}\approx{\mathbf{g}_{\phi}}^{-1}$ and train the inverse map with a constraint loss to ensure $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})$ is close to the constraint set for all $\tilde{\mathbf{x}}$ that we are interested in (we define this formally in Section 3.1). An MDM can be trained on the pushforward of the data distribution through $\mathbf{g}_{\phi}$, and its generated samples can be mapped to the constraint set via $\mathbf{f}_{\psi}$. Although not inherently restricted to diffusion models, our approach maintains the many advantages of diffusion models, including expressive generation, simulation-free training (Song et al., 2021b), and tractable computation of probability densities (Liu et al., 2024). One can also adapt existing diffusion-based inverse solvers for the mirror space and enforce constraints with the inverse map.
@@
Our experiments show improved constraint satisfaction for various physics-based, geometric, and semantic constraints. We also discuss ablation studies and adapt a popular diffusion-based inverse solver to solve constrained inverse problems, in particular data assimilation with PDE constraints.
@@ 2.1 Constrained generative models
Explicitly incorporating a known constraint into a generative model poses benefits such as data efficiency (Ganchev et al., 2010; Batzner et al., 2022), generalization capabilities (Köhler et al., 2020), and feasibility of samples (Giannone et al., 2023). Some methods leverage equivariant neural networks (Satorras et al., 2021; Geiger & Smidt, 2022; Thomas et al., 2018) for symmetry (Allingham et al., 2022; Niu et al., 2020; Klein et al., 2024; Song et al., 2024; Boyda et al., 2021; Rezende et al., 2019; Garcia Satorras et al., 2021; Köhler et al., 2020; Midgley et al., 2024; Dey et al., 2020; Xu et al., 2022; Hoogeboom et al., 2022; Yim et al., 2023) but do not generalize to other types of constraints or generative models (Klein et al., 2024; Song et al., 2024; Boyda et al., 2021; Rezende et al., 2019; Garcia Satorras et al., 2021; Köhler et al., 2020; Midgley et al., 2024; Niu et al., 2020; Xu et al., 2022; Hoogeboom et al., 2022; Yim et al., 2023; Dey et al., 2020; Allingham et al., 2022). Previous methods for constrained diffusion models (De Bortoli et al., 2022; Huang et al., 2022; Fishman et al., 2023a; Lou & Ermon, 2023; Liu et al., 2024) make strong assumptions about the constraint, such as being characterized as a Riemannian manifold (De Bortoli et al., 2022; Huang et al., 2022), having a well-defined reflection operator (Lou & Ermon, 2023) or projection operator (Christopher et al., 2024), or corresponding to a convex constraint set (Fishman et al., 2023a; Liu et al., 2024). Fishman et al. (2023b) proposed a diffusion model that incorporates Metropolis-Hastings steps to work with general constraints, but impractically high rejection rates may occur with constraints that are challenging to satisfy, such as a PDE constraint.
An alternative approach is to introduce a soft constraint penalty when training the generative model (Ganchev et al., 2010; Mann & McCallum, 2007; Chang et al., 2007; Giannone et al., 2023; Daras et al., 2024; Upadhyay et al., 2023). However, evaluating the constraint loss of generated samples during training may be prohibitively expensive. Instead, one could add constraint-violating training examples (Giannone et al., 2023), but it is often difficult to procure useful negative examples. In contrast, our approach does not alter the training objective of the generative model.
@@ 2.2 Mirror maps
For any convex constraint set $\mathcal{C}\subseteq\mathbb{R}^{d}$, one can define a mirror map that maps from $\mathcal{C}$ to $\mathbb{R}^{d}$. This is done by defining a mirror potential $\phi:\mathcal{C}\to\mathbb{R}$ that is continuously-differentiable and strongly-convex (Bubeck et al., 2015; Tan et al., 2023). The mirror map is the gradient $\nabla\phi:\mathcal{C}\to\mathbb{R}^{d}$ (Liu et al., 2024). Every mirror map has an inverse $(\nabla\phi)^{-1}:\mathbb{R}^{d}\to\mathcal{C}$, which, unlike the forward mirror map, is not necessarily the gradient of a strongly-convex function (Zhou, 2018; Tan et al., 2023). Mirror maps have been used for constrained optimization (Beck & Teboulle, 2003) and sampling (Hsieh et al., 2018; Li et al., 2022; Liu et al., 2024). Although true mirror maps exist only for convex constraints, we seek to generalize the concept to learn approximate mirror maps to handle non-convex constraints. Recent work suggested learned mirror maps for convex optimization (Tan et al., 2023) and reinforcement learning (Alfano et al., 2024) but did not tackle constrained generative modeling. Our work proposes a novel training objective for learning mirror maps with the goal of constraining generative models to satisfy arbitrary constraints.
@@ 2.3 Diffusion models
A diffusion model learns to generate new data samples through a gradual denoising process (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019; Song et al., 2021b; Kingma et al., 2021). The diffusion, or noising, process can be modeled as a stochastic differential equation (SDE) (Song et al., 2021b) that induces a time-dependent distribution $p_{t}$, where $p_{0}=p_{\text{data}}$ (the target data distribution) and $p_{T}\approx\mathcal{N}(\mathbf{0},\mathbf{I})$. The diffusion model learns to reverse the denoising process by modeling the score function of $p_{t}$, defined as $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. In the context of imaging problems, the score function is often parameterized using a convolutional neural network (CNN) with parameters $\theta$, which we denote by $\mathbf{s}_{\theta}$. The score model is trained with a denoising-based objective (Hyvärinen & Dayan, 2005; Vincent, 2011; Song et al., 2021b) that allows for simulation-free training (i.e., simulating the forward noising process is not necessary during training). The score function appears in a reverse-time SDE (Song et al., 2021b; Anderson, 1982) that can be used to sample from the clean image distribution $p_{0}$ by first sampling a noise image from $\mathcal{N}(\mathbf{0},\mathbf{I})$ and then gradually denoising it. Our work addresses the problem that diffusion models are often not sensitive to visually-subtle constraints on the data. Our proposed NAMM allows for a diffusion model to be trained in an unconstrained space yet satisfy the desired constraint by construction.
@@ 3 Method
We now describe neural approximate mirror maps (NAMMs). We focus on diffusion models, but any generative model can be trained in the learned mirror space (see Appendix A for results with a VAE). We denote the constrained image distribution by $p_{\text{data}}$ and the (not necessarily convex) constraint set by $\mathcal{M}\subseteq\mathbb{R}^{d}$. Images in the constrained and mirror spaces are denoted by $\mathbf{x}$ and $\tilde{\mathbf{x}}$, respectively. The pushforward of the data distribution $p_{\text{data}}$ through a mirror map $\mathbf{g}_{\phi}$ is denoted by ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$.
@@ 3.1 Learning the forward and inverse mirror maps
Figure 2: NAMM training illustration. Given data that lie on a constraint manifold $\mathcal{M}$ (e.g., the hyperplane of images with the same total brightness), we jointly train an approximate mirror map $\mathbf{g}_{\phi}$ and its approximate inverse $\mathbf{f}_{\psi}$. After mapping data $\mathbf{x}\sim p_{\text{data}}$ to the mirror space as $\mathbf{g}_{\phi}(\mathbf{x})$, we perturb them with additive Gaussian noise whose standard deviation can be anywhere between 00 and $\sigma_{\text{max}}$. The inverse map $\mathbf{f}_{\psi}$ is trained to map these perturbed samples back onto $\mathcal{M}$.
@@
A true inverse mirror map satisfies cycle consistency and constraint satisfaction on all of $\mathbb{R}^{d}$, so ideally $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})=\mathbf{g}_{\phi}^{-1}(\tilde{\mathbf{x}})$ and $\mathbf{f}_{\psi}(\tilde{\mathbf{x}})\in\mathcal{M}$ for all $\tilde{\mathbf{x}}\in\mathbb{R}^{d}$. But since it would be computationally infeasible to optimize $\mathbf{f}_{\psi}$ over all possible points in $\mathbb{R}^{d}$, we instead optimize it over distributions that we would expect the inverse map to face in practice in the context of generative models. That is, we only need $\mathbf{f}_{\psi}$ to be valid for samples from an MDM trained on ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$, which we refer to as the mirror distribution. To make $\mathbf{f}_{\psi}$ robust to learning/sampling error of the MDM, we consider a sequence of noisy distributions in the mirror space, each corresponding to adding Gaussian noise to samples from ${(\mathbf{g}_{\phi})}_{\sharp}p_{\text{data}}$, which, for a maximum perturbation level $\sigma_{\text{max}}$, we denote by
@@
We train $\mathbf{f}_{\psi}$ to be a valid inverse mirror map only for points from these noisy mirror distributions. Since we do not know a priori how much noise the MDM samples will contain, we consider all possible noise levels up to $\sigma_{\text{max}}$ for robustness (see Appendix D.2 for an ablation study of this choice).
We define a cycle-consistency loss (Zhu et al., 2017) that covers the forward and inverse directions and evaluates the inverse direction for the entire sequence of distributions defined in Equation 2:
@@
Let $\ell_{\text{constr}}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be a differentiable constraint distance that measures the distance from an input image to the constraint set. We define the following constraint loss to encourage $\mathbf{f}_{\psi}$ to map points from the noisy mirror distributions (Equation 2) to the constraint set:
@@
We use Monte-Carlo to approximate the expectations in the objective over the noisy mirror distributions with $\sigma\sim\mathcal{U}([0,\sigma_{\text{max}}])$ and approximately solve Equation 1 with stochastic gradient descent.
@@
We parameterize $\mathbf{g}_{\phi}$ as the gradient of an input-convex neural network (ICNN) following the implementation of Tan et al. (2023). For convex constraints, this satisfies the theoretical requirement that $\mathbf{g}_{\phi}$ be the gradient of a strongly-convex function. Even for non-convex constraints, this choice brings practical benefits, as we discuss in Section 4.3. We note that $\mathbf{g}_{\phi}$ is not a true mirror map since $\mathcal{M}$ is not assumed to be convex, and $\mathbf{g}_{\phi}$ is defined on all of $\mathbb{R}^{d}$ instead of just on $\mathcal{M}$. We parameterize $\mathbf{f}_{\psi}$ as a ResNet-based CNN similar to the one used in CycleGAN (Zhu et al., 2017).